Find the area of a triangle if the base is 12x^7, and the height is 5x^8. 2nd problem- Find the area of a rectangle if the length is 36y^3 and the width is 14y^5
Answer by checkley77(12844)

Area=bh/2=12x^7*5x^6=60x^13 ans.
Area=LW=36y^3*14y^5=504y^8 ans.